OpenStudy (anonymous):

Pyramid A is a square pyramid with a base side length of 12 inches and a height of 8 inches. Pyramid B is a square pyramid with a base side length of 24 inches and a height of 16 inches. How many times bigger is the volume of pyramid B than pyramid A? Numerical Answers Expected!

OpenStudy (anonymous):

@kat_here2learn

OpenStudy (kat_here2learn):

Area of the base * Height * 1/3 V(A) = 12*8*1/3 V(A)=31.97 V(B)=24*16*1/3 V(B)=127.9V(A) = 12^2*8*1/3 V(A)=383.96 V(B) = 24^2*16*1/3 V(B) = 3,071.97 3,071.97/383.96= 8

OpenStudy (anonymous):

can u help with a few more?

OpenStudy (kat_here2learn):

V(A) is 8 times smaller than v(b) Though I suggest you round answers to 384 and 3072

OpenStudy (kat_here2learn):

i can try!

OpenStudy (anonymous):

so the answer is 8 right?

OpenStudy (kat_here2learn):

yes sir

OpenStudy (anonymous):

its crazy how u did that so fast

OpenStudy (kat_here2learn):

math is like second nature to me these are quiet simple questions

OpenStudy (anonymous):

oml really?

OpenStudy (kat_here2learn):

are you confused would you like me to explain the anser further?

OpenStudy (kat_here2learn):

need more help with any other questions?

OpenStudy (anonymous):

for now no if do i will tag you

OpenStudy (kat_here2learn):

okay have a nice day!

OpenStudy (anonymous):

but no i wish these were simple too me lol